# On the Convergence of Muon and Beyond

### v2

@@ Abstract
The Muon optimizer has demonstrated remarkable empirical success in handling matrix-structured parameters for training neural networks. However, a significant gap persists between its practical performance and theoretical understanding. Existing analyses indicate that the standard Muon variant achieves only a suboptimal convergence rate of $\mathcal{O}(T^{-1/4})$ in stochastic non-convex settings, where $T$ denotes the number of iterations. To explore the theoretical limits of the Muon framework, we develop and analyze two momentum-based variance-reduced variants: a one-batch version (Muon-MVR1) and a two-batch version (Muon-MVR2). We provide the first rigorous proof that incorporating a variance-reduction mechanism enables Muon-MVR2 to attain an optimal convergence rate of $\tilde{\mathcal{O}}(T^{-1/3})$, thereby matching the theoretical lower bound for this class of problems. Moreover, our analysis establishes convergence guarantees for Muon variants under the Polyak-Łojasiewicz (PŁ) condition. Extensive experiments on vision (CIFAR-10) and language (C4) benchmarks corroborate our theoretical findings on per-iteration convergence. Overall, this work provides the first proof of optimality for a Muon-style optimizer and clarifies the path toward developing more practically efficient, accelerated variants.
@@ 1 Introduction
The immense computational cost of pre-training Large Language Models (LLMs) has spurred a surge of research into novel optimization methods designed to enhance parameter efficiency and training stability Hoffmann et al. (2022); Liu et al. (2023); Chen et al. (2023); Vyas et al. (2025); Pethick et al. (2025); Yuan et al. (2024). Among these, methods based on matrix orthogonalization have recently garnered significant attention from both academia and industry Jordan et al. (2024); Liu et al. (2025a). In particular, the Muon optimizer has emerged as a notable milestone due to its impressive empirical performance Liu et al. (2025a); An et al. (2025); Liu et al. (2025b); Shah et al. (2025). However, despite its practical success, the theoretical understanding of Muon’s underlying mechanisms has surprisingly lagged behind, with existing convergence analyses being fraught with limitations and even critical fallacies.
Specifically, the current theoretical exploration of Muon’s convergence faces three primary obstacles. First, some analyses are predicated on flawed mathematical assumptions, such as the erroneous use of the inequality $\|\mathbf{S}^{-1}\|_{2}\leq 1/\|\mathbf{S}\|_{2}$ in a key part of their proof, which casts serious doubt on the validity of their convergence claims Li & Hong (2025). Second, existing analyses have not established convergence to a stationary point without reliance on problem dimension or batch size; the available bounds become valid only when the batch size is sufficiently large, so the basic requirement for a trustworthy optimizer remains unfulfilled Sato et al. (2025); Sfyraki & Wang (2025). Finally, the most rigorous existing work analyzes both the standard Muon algorithm (like Algorithm 1, Option MVR1 with $\gamma=0$) and its simplified variant , considering their convergence to a non-standard $\epsilon$-nuclear norm stationary point. This leaves the behavior of the Nesterov-Accelerated Muon algorithm (like Algorithm 1, Option MVR1), as discussed in Liu et al. (2025a); Sato et al. (2025), and Variance-Reduction Muon (like Algorithm 1, Option MVR2) under standard settings as an open question Shen et al. (2025).
@@
  • Regarding the optimal convergence rate, we are the first to demonstrate that the Variance-Reduction Muon-MVR2 algorithm (Algorithm 1, Option MVR2) can overcome the bottleneck of traditional momentum methods, achieving the optimal ergodic convergence rate of $\tilde{\mathcal{O}}(T^{-1/3})$.
  • Under the Polyak-Łojasiewicz (PL) condition, we prove that our proposed algorithms, Muon-MVR1, and Muon-MVR2, all achieve sublinear convergence rates. More precisely, we demonstrate that Muon-MVR1 converge at a rate of $\tilde{\mathcal{O}}(1/T^{1/2})$, and Muon-MVR2 converges at an accelerated rate of $\tilde{\mathcal{O}}(1/T^{2/3})$
We also refer readers to see the main contributions of our works and comparisons with the existing works in Table 1.
Table 1: Comparison for existing Muon-type analyses with ours
Smooth
@@
Convergence
Rate
Li & Hong (2025) $L$ Option MVR1($\gamma=0$) $\mathcal{O}(1)$ $\mathcal{O}(T^{-1/4})$
Sato et al. (2025) $L$ Option MVR1($\gamma=0$) $\mathcal{O}(\epsilon^{-1})$ $\mathcal{O}(n)$
Sato et al. (2025) $L$ Option MVR1 $\mathcal{O}(\epsilon^{-1})$ $\mathcal{O}(n)$
Sfyraki & Wang (2025) $L_{+}$ Option MVR2 $\mathcal{O}(\epsilon^{-1})$ $\mathcal{O}(T^{-1/3})$
Shen et al. (2025) $L$ Option MVR1($\gamma=0$) $\mathcal{O}(1)$ $\mathcal{O}(T^{-1/4})$
Ours $L$ Option MVR1($\gamma=0$) $\mathcal{O}(1)$ $\tilde{\mathcal{O}}(T^{-1/4})$ Th4.3
Ours $L$ Option MVR1 $\mathcal{O}(1)$ $\tilde{\mathcal{O}}(T^{-1/4})$ Th4.3
Ours $L_{+}$ Option MVR2 $\mathcal{O}(1)$ $\tilde{\mathcal{O}}(T^{-1/3})$ Th4.4
@@
  • b
    Although increasing the batch size can mitigate the impact of stochastic noise Sato et al. (2025), these methods still fail to converge to a stationary point and cannot eliminate the influence of dimensionality.
@@
    The results from Sfyraki & Wang (2025) on Option MVR2 are the closest to ours. However, their method requires a large initial batch of size $\mathcal{O}(\epsilon^{-1})$, although the batch size can be reduced to 1 in subsequent iterations.
2 Preliminaries and Problem Setup
2.1 Notations
We denote scalars by non-bold letters (e.g., $a,A$), vectors in $\mathbb{R}^{d}$ by bold lowercase letters (e.g., $\mathbf{a}$), and matrices by bold uppercase letters (e.g., $\mathbf{A}$). The space $\mathbb{R}^{d}$ is endowed with the Euclidean inner product $\langle\mathbf{x},\mathbf{y}\rangle_{2}:=\mathbf{x}^{\top}\mathbf{y}$ and norm $\|\mathbf{x}\|_{2}$. For matrices, we employ the Frobenius inner product $\langle\mathbf{A},\mathbf{B}\rangle_{F}:=\text{tr}(\mathbf{A}^{\top}\mathbf{B})$ and the corresponding norm $\|\mathbf{A}\|_{F}$. The nuclear norm, denoted by $\|\mathbf{A}\|_{*}$, is defined as the sum of the singular values of the matrix, $\|\mathbf{A}\|_{*}=\sum_{i}\sigma_{i}(\mathbf{A})$. Throughout the paper, $[m]$ denotes the set of integers $\{1,2,\ldots,m\}$, and $\mathbb{N}$ denotes the set of non-negative integers.
The model is parameterized by a matrix $\mathbf{X}\in\mathbb{R}^{m\times n}$. Without loss of generality, we assume $m\geq n$, so the rank of the matrix is at most $n$. The model is optimized by minimizing the empirical loss function $f(\mathbf{X}):=\frac{1}{N}\sum_{i\in[N]}f_{i}(\mathbf{X})$, where $N$ is the number of training data points, and $f_{i}(\mathbf{X})$ is the loss function for $\mathbf{X}\in\mathbb{R}^{m\times n}$ with respect to the $i$-th training data point $\mathbf{z}_{i}$ (for $i\in[N]$). Let $\xi$ be a random variable that is independent of $\mathbf{X}\in\mathbb{R}^{m\times n}$, and let $\mathbb{E}_{\xi}[\mathbf{X}]$ denote the expectation of a random variable $\mathbf{X}$ with respect to $\xi$.
2.2 Problem Setup
We consider minimizing the objective function as follows
$\min_{\mathbf{X}}f(\mathbf{X}),\mathrm{~where~}f(\mathbf{X})=\mathbb{E}_{\xi\sim\mathcal{D}}[f(\mathbf{X};\xi)].$ (1)
For problem (1), let $f:\mathbb{R}^{m\times n}\rightarrow\mathbb{R}$ be a function where $\mathbf{X}\in\mathbb{R}^{m\times n}$ and $\xi$ represents a random variable, such as a training data point, sampled from an unknown data distribution $\mathcal{D}$. We assume that $f$ is differentiable and possibly nonconvex.
@@ 3 Revisiting the Muon Algorithms
1:Require Initial parameters $\mathbf{X}_{0}\in\mathbb{R}^{m\times n}$, learning rate $\eta_{t}>0$, momentum parameter $\beta_{t}\in[0,1)$, variance-reduction parameter $\gamma\in[0,1]$, initial momentum $\mathbf{M}_{0}=0$, $\nabla f(\mathbf{X}_{0};\xi)=0$.
@@
11:Return Final parameters $\mathbf{X}_{t}$
In practice, Algorithm 1 leverages a quintic Newton–Schulz iteration to efficiently compute a rank-preserving normalization of any input matrix $\mathbf{M}_{t}\in\mathbb{R}^{m\times n}$. This operation can be precisely defined via the Compact Singular Value Decomposition (SVD) of $\mathbf{M}_{t}$, expressed as $\mathbf{M}_{t}=\mathbf{U}_{r}\mathbf{\Sigma}_{r}\mathbf{V}_{r}^{\top}$, where $r$ is the rank of $\mathbf{M}_{t}$ and $\mathbf{\Sigma}_{r}\in\mathbb{R}^{r\times r}$ contains only the $r$ positive singular values. The operator computes the matrix $\mathbf{O}_{t}\in\mathbb{R}^{m\times n}$ given by:
$\mathbf{O}_{t}=\mathbf{U}_{r}\mathbf{V}_{r}^{\top}$
This is equivalent to setting all positive singular values of $\mathbf{M}_{t}$ to 1 while leaving its null space unchanged, thus preserving the original rank $r$. It is worth noting that this formulation is algebraically equivalent to using the Moore-Penrose pseudoinverse: $((\mathbf{M}_{t}\mathbf{M}_{t}^{\top})^{+})^{1/2}\mathbf{M}_{t}$.
Only in the specific case where $\mathbf{M}_{t}$ is full-rank does this result coincide with the solution to the Orthogonal Procrustes problem, $\min_{\mathbf{O}}\|\mathbf{O}-\mathbf{M}_{t}\|_{F}^{2}$ s.t. $\mathbf{O}^{\top}\mathbf{O}=\mathbf{I}$. In the subsequent theoretical analysis, we assume that this operator can be solved exactly. Building upon this core operator, we investigate the following update rules:
@@
$\mathbf{M}_{t}=\beta_{t}\mathbf{M}_{t-1}+(1-\beta_{t})\nabla f(\mathbf{X}_{t};\xi_{t})+\gamma\cdot\beta_{t}\cdot(\nabla f(\mathbf{X}_{t};\xi_{t})-\nabla f(\mathbf{X}_{t-1};\xi_{t-1}))$ (2)
Further analysis reveals that when the parameters are set as $\beta_{t}=\mu$ and $\gamma=1-\beta_{t}$, the iterative process of Algorithm 1, subject to a rescaling of $\mathbf{M}_{t}$, becomes mathematically equivalent to the update scheme (3) of the Muon algorithm Jordan et al. (2024); Liu et al. (2025a). The Muon update is given by: $\mathbf{M}_{t}=\mu\mathbf{M}_{t-1}+\nabla f(\mathbf{X}_{t};\xi_{t})+\mu(\nabla f(\mathbf{X}_{t};\xi_{t})-\nabla f(\mathbf{X}_{t-1};\xi_{t-1}))$
$\mathbf{C}_{t}$ $=\mu\mathbf{C}_{t-1}+\nabla f(\mathbf{X}_{t};\xi_{t})$ (3)
$\mathbf{M}_{t}$ $=\mu\mathbf{C}_{t}+\nabla f(\mathbf{X}_{t};\xi_{t})$
$\mathbf{O}_{t}$ $=\text{Newton-Schulz}(\mathbf{M}_{t})$
$\mathbf{X}_{t+1}$ $=\mathbf{X}_{t}-\eta\mathbf{O}_{t}$
Muon-MVR1 reduces to the EMA stochastic gradient estimator Kingma & Ba (2014) update Rule (4) as a special case when $\gamma=0$, and it can also be regarded as an approximation of MVR2.
$\mathbf{M}_{t}=\beta_{t}\mathbf{M}_{t-1}+(1-\beta_{t})\nabla f(\mathbf{X}_{t};\xi_{t})$ (4)
@@
$\mathbf{M}_{t}=\beta_{t}\mathbf{M}_{t-1}+(1-\beta_{t})\nabla f(\mathbf{X}_{t};\xi_{t})+\gamma\cdot\beta_{t}\cdot(\nabla f(\mathbf{X}_{t};\xi_{t})-\nabla f(\mathbf{X}_{t-1};\xi_{t}))$ (5)
The key distinction between this rule and Rule (2) lies in the term subtracted from the current stochastic gradient, $\nabla f(\mathbf{X}_{t};\xi_{t})$. Specifically, Rule (5) subtracts the previous gradient computed using the current data sample $\xi_{t}$, namely $\nabla f(\mathbf{X}_{t-1};\xi_{t})$, a crucial modification designed to cancel the variance introduced by $\xi_{t}$ Yuan et al. (2024). In contrast, Rule (2) reuses the gradient based on the previous data sample $\xi_{t-1}$ (i.e., $\nabla f(\mathbf{X}_{t-1};\xi_{t-1})$), effectively acting as a computationally efficient approximation of Rule (5).
This approximation is justified because the difference between the two approaches is a noise term, $\Delta^{\text{Noise}}_{t-1}=\nabla f(\mathbf{X}_{t-1};\xi_{t})-\nabla f(\mathbf{X}_{t-1};\xi_{t-1})$. Under the standard assumption that stochastic gradients have a bounded variance, $\mathbb{E}_{\xi}[\|\nabla f(\mathbf{X};\xi)-\nabla f(\mathbf{X})\|_{F}^{2}]\leq\sigma^{2}$, the variance of this noise term is well-controlled and satisfies $\mathbb{E}[\|\Delta^{\text{Noise}}_{t-1}\|_{F}^{2}]\leq 2\sigma^{2}$. While Rule (2) offers the practical benefit of requiring only one gradient evaluation per step versus two for Rule (5), this structural difference leads to fundamentally different convergence properties. Our theoretical analysis demonstrates that the Muon-variant algorithm employing the stricter variance-reduction mechanism of Rule (5) achieves a convergence rate of $\tilde{\mathcal{O}}(T^{-1/3})$ in Theorem 4.2.

### v3

@@ Abstract
The Muon optimizer has demonstrated remarkable empirical success in handling matrix-structured parameters for training neural networks. However, a significant gap remains between its practical performance and theoretical understanding. Existing analyses show that the Muon variants achieve only a suboptimal iteration complexity of $\mathcal{O}(T^{-1/4})$ in stochastic non-convex settings, where $T$ denotes the number of iterations. To explore the theoretical limits of the Muon framework, we analyze two Momentum-based Variance-Reduced variants: a one-batch version (Muon-MVR1) and a two-batch version (Muon-MVR2). We provide the first rigorous proof that incorporating variance reduction enables Muon-MVR2 to attain the optimal iteration complexity of $\tilde{\mathcal{O}}(T^{-1/3})$, thereby matching the theoretical lower bound for this class of problems. Furthermore, our analysis establishes last-iterate convergence guarantees for Muon variants under the Polyak-Łojasiewicz (PŁ) condition. Extensive experiments on vision (CIFAR-10) and language (C4) benchmarks corroborate our theoretical findings on per-iteration convergence. Overall, this work offers the first proof of optimality for a Muon-style optimizer and clarifies the path toward developing more practically efficient, accelerated variants.
@@ 1 Introduction
The immense computational cost of pre-training Large Language Models (LLMs) has spurred a surge of research into novel optimization methods designed to enhance parameter efficiency and training stability hoffmann2022training; Liu2023SophiaAS; Chen2023SymbolicDO; vyas2025soap; pethick2025training; Yuan2024MARSUT. Among these, methods based on matrix orthogonalization have recently garnered significant attention from both academia and industry jordan2024muon; Liu2025MuonIS. In particular, the Muon optimizer has emerged as a notable milestone due to its impressive empirical performance Liu2025MuonIS; An2025ASGOAS; liu2025cosmos; shah2025practical. However, despite its practical success, the theoretical understanding of Muon’s underlying mechanisms has surprisingly lagged behind, with existing convergence analyses being fraught with limitations and even critical fallacies.
Specifically, the current theoretical exploration of Muon’s convergence faces three primary obstacles. First, existing analyses have not established convergence to a stationary point without reliance on problem dimension or batch size; the available bounds become valid only when the batch size is sufficiently large, so the basic requirement for a trustworthy optimizer remains unfulfilled Sato2025ConvergenceBA; sfyraki2025lions. Second, some analyses are predicated on flawed mathematical assumptions, such as the erroneous use of the inequality $\|\mathbf{S}^{-1}\|_{2}\leq 1/\|\mathbf{S}\|_{2}$ in a key part of their proof, which casts serious doubt on the validity of their convergence claims Li2025ANO. Finally, the most rigorous existing work analyzes both the standard Muon algorithm (like Algorithm 1, Option MVR1 with $\gamma=0$) and its simplified variant , considering their convergence to a non-standard $\epsilon$-nuclear norm stationary point. This leaves the behavior of the Nesterov-Accelerated Muon algorithm (like Algorithm 1, Option MVR1), as discussed in Liu2025MuonIS; Sato2025ConvergenceBA, and Variance-Reduction Muon (like Algorithm 1, Option MVR2) under standard settings as an open question shen2025convergence.
@@
  • Regarding the optimal iteration complexity, we are the first to demonstrate that the Variance-Reduction Muon-MVR2 algorithm (Algorithm 1, Option MVR2) can overcome the bottleneck of traditional momentum methods, achieving the optimal iteration complexity of $\tilde{\mathcal{O}}(T^{-1/3})$.
  • Under the Polyak-Łojasiewicz (PL) condition, we prove that our proposed algorithms, Muon-MVR1, and Muon-MVR2, all achieve sublinear convergence rates. More precisely, we demonstrate that Muon-MVR1 converge at a rate of $\tilde{\mathcal{O}}(T^{-1/2})$, and Muon-MVR2 converges at an accelerated rate of $\tilde{\mathcal{O}}(T^{-2/3})$
Table 1 summarizes the main contributions of our work and compares them with existing methods.
Smootha
@@
Last-iterate
Conv. Rate
Li2025ANO $L$ MVR1($\gamma=0$) $\mathcal{O}(1)$ $\mathcal{O}(T^{-1/4})$
Sato2025ConvergenceBA $L$ MVR1($\gamma=0$) $\mathcal{O}(\epsilon^{-1})$ $\mathcal{O}(T^{-1})+\mathcal{O}(1)^{c}$
Sato2025ConvergenceBA $L$ MVR1 $\mathcal{O}(\epsilon^{-1})$ $\mathcal{O}(T^{-1})+\mathcal{O}(1)^{c}$
sfyraki2025lions $L_{+}$ MVR2 $\mathcal{O}(\epsilon^{-1})^{d}$ $\mathcal{O}(T^{-1/3})$
shen2025convergence $L$ MVR1($\gamma=0$) $\mathcal{O}(1)$ $\mathcal{O}(T^{-1/4})$
Ours $L$ MVR1($\gamma=0$) $\mathcal{O}(1)$ $\tilde{\mathcal{O}}(T^{-1/4})$ $\tilde{\mathcal{O}}(T^{-1/2})$
Ours $L$ MVR1 $\mathcal{O}(1)$ $\tilde{\mathcal{O}}(T^{-1/4})$ $\tilde{\mathcal{O}}(T^{-1/2})$
Ours $L_{+}$ MVR2 $\mathcal{O}(1)$ $\tilde{\mathcal{O}}(T^{-1/3})$ $\tilde{\mathcal{O}}(T^{-2/3})$
@@
    Although increasing the batch size can mitigate the impact of stochastic noise Sato2025ConvergenceBA, these methods still fail to converge to a stationary point and cannot eliminate the influence of dimensionality.
  • d
    The results from sfyraki2025lions on Option MVR2 are the closest to ours. However, their method requires a large initial batch of size $\mathcal{O}(\epsilon^{-1})$, although the batch size can be reduced to 1 in subsequent iterations.
Table 1: Comparison of Existing Muon-type Analyses with Ours.
$\bullet$ Organizations. The rest of the paper is organized as follows. Section 2 reviews existing variants of the Muon algorithm. Section 3 addresses iteration complexity and last-iterate convergence rate. Section 4 presents experimental results demonstrating the effectiveness of our method. Related work is reviewed in Section 5, and conclusions are drawn in Section 6.
$\bullet$ Notations. We denote scalars by non-bold letters (e.g., $a,A$), vectors in $\mathbb{R}^{d}$ by bold lowercase letters (e.g., $\mathbf{a}$), and matrices by bold uppercase letters (e.g., $\mathbf{A}$). The space $\mathbb{R}^{d}$ is endowed with the Euclidean inner product $\langle\mathbf{x},\mathbf{y}\rangle_{2}:=\mathbf{x}^{\top}\mathbf{y}$ and norm $\|\mathbf{x}\|_{2}$. For matrices, we employ the Frobenius inner product $\langle\mathbf{A},\mathbf{B}\rangle_{F}:=\text{tr}(\mathbf{A}^{\top}\mathbf{B})$ and the corresponding norm $\|\mathbf{A}\|_{F}$. The nuclear norm, denoted by $\|\mathbf{A}\|_{*}$, is defined as the sum of the singular values of the matrix, $\|\mathbf{A}\|_{*}=\sum_{i}\sigma_{i}(\mathbf{A})$. Throughout the paper, $[m]$ denotes the set of integers $\{1,2,\ldots,m\}$, and $\mathbb{N}$ denotes the set of non-negative integers. The model is parameterized by a matrix $\mathbf{X}\in\mathbb{R}^{m\times n}$. Without loss of generality, we assume $m\geq n$, so the rank of the matrix is at most $n$. The model is optimized by minimizing the empirical loss function $f(\mathbf{X}):=\frac{1}{N}\sum_{i\in[N]}f_{i}(\mathbf{X})$, where $N$ is the number of training data points, and $f_{i}(\mathbf{X})$ is the loss function for $\mathbf{X}\in\mathbb{R}^{m\times n}$ with respect to the $i$-th training data point $\mathbf{z}_{i}$ (for $i\in[N]$). Let $\xi$ be a random variable that is independent of $\mathbf{X}\in\mathbb{R}^{m\times n}$, and let $\mathbb{E}_{\xi}[\mathbf{X}]$ denote the expectation of a random variable $\mathbf{X}$ with respect to $\xi$.
@@ 2 Revisiting the Muon Algorithms
We consider the following optimization problem:
$\min_{\mathbf{X}\in\mathbb{R}^{m\times n}}f(\mathbf{X}),\mathrm{~where~}f(\mathbf{X})=\mathbb{E}_{\xi\sim\mathcal{D}}[f(\mathbf{X};\xi)],$ (1)
where $f:\mathbb{R}^{m\times n}\rightarrow\mathbb{R}$ is the loss function, $\mathbf{X}$ denotes the decision variable, and $\xi$ represents a random variable(e.g., a training data sample) drawn from an unknown distribution $\mathcal{D}$. We assume that $f$ is differentiable and possibly nonconvex.
The Muon optimizer begins by computing a momentum-based variance-reduced gradient update, similar in spirit to ADAM Kingma2014AdamAM, STORM Cutkosky2019MomentumBasedVR, and SGD with Nesterov momentum sutskever2013importance. The momentum term $\mathbf{M}_{t}$ is then projected onto the set of orthogonal matrices. This orthogonalization step equalizes the singular values, ensuring that no principal component direction dominates the optimization. Finally, the resulting scaled orthogonal matrix is used to update the model parameters. The Muon algorithm is summarized in Algorithm 1.
@@
1:Input: Initial parameters $\mathbf{X}_{0}\in\mathbb{R}^{m\times n}$, learning rate $\eta_{t}>0$, momentum parameter $\beta_{t}\in[0,1)$, variance-reduction parameter $\gamma\in[0,1]$, initial momentum $\mathbf{M}_{0}=0$, $\nabla f(\mathbf{X}_{0};\xi)=0$.
@@
11:Output: Final parameters $\mathbf{X}_{t}$
This orthogonalization step can be obtained from the Singular Value Decomposition (SVD) of $\mathbf{M}_{t}$:
$\mathbf{M}_{t}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{\top},\qquad\mathbf{O}_{t}=\mathbf{U}\mathbf{V}^{\top}.$
Equivalently, we have $\mathbf{O}_{t}=(\mathbf{M}_{t}\mathbf{M}_{t}^{\top})^{-1/2}\mathbf{M}_{t}$, which shows that orthogonalization reduces to computing an inverse square root rather than a full SVD. Since forming and decomposing $\mathbf{M}_{t}\mathbf{M}_{t}^{\top}$ remains costly, the popular Muon implementation jordan2024muon uses the quintic Newton–Schulz iteration to approximate the inverse square root. This recurrence converges in only a few steps (typically five), producing a numerically stable, rank-preserving orthogonalization of $\mathbf{M}_{t}$ that is nearly as accurate as SVD but far more efficient.
Algorithm 1 incorporates two distinct strategies for momentum-based variance reduction, termed Muon-MVR1 and Muon-MVR2. These options present a fundamental trade-off between computational efficiency and theoretical rigor. MVR2 implements a principled variance reduction scheme at the cost of two gradient evaluations per step, while MVR1 serves as a computationally cheaper, single-gradient approximation. We detail both below.
@@
$\mathbf{M}_{t}=\beta_{t}\mathbf{M}_{t-1}+(1-\beta_{t})\nabla f(\mathbf{X}_{t};\xi_{t})+\gamma\cdot\beta_{t}\cdot(\nabla f(\mathbf{X}_{t};\xi_{t})-\nabla f(\mathbf{X}_{t-1};\xi_{t-1})).$ (2)
The primary advantage of this formulation is its computational efficiency, as it requires only one stochastic gradient evaluation per iteration. This update rule is flexible:
(i) When $\gamma=0$, Rule (2) degenerates to the standard exponential moving average (EMA) of gradients (Rule (3)), a stochastic gradient estimator widely used in optimizers like Adam Kingma2014AdamAM.
$\mathbf{M}_{t}=\beta_{t}\mathbf{M}_{t-1}+(1-\beta_{t})\nabla f(\mathbf{X}_{t};\xi_{t}).$ (3)
(ii) With the specific parameterization ($\beta_{t}=\mu,\gamma=1-\beta_{t}$), it can be shown that Rule (2) becomes mathematically equivalent to the update scheme of the Muon optimizer jordan2024muon; Liu2025MuonIS, given by:
$\mathbf{C}_{t}$ $=\mu\mathbf{C}_{t-1}+\nabla f(\mathbf{X}_{t};\xi_{t}),$ (4)
$\mathbf{M}_{t}$ $=\mu\mathbf{C}_{t}+\nabla f(\mathbf{X}_{t};\xi_{t}).$
@@
$\mathbf{M}_{t}=\beta_{t}\mathbf{M}_{t-1}+(1-\beta_{t})\nabla f(\mathbf{X}_{t};\xi_{t})+\gamma\cdot\beta_{t}\cdot(\nabla f(\mathbf{X}_{t};\xi_{t})-\nabla f(\mathbf{X}_{t-1};\xi_{t})).$ (5)
The key distinction from MVR1 is the correction term. MVR2 subtracts the gradient computed on the previous parameters but with the current mini-batch, i.e., $\nabla f(\mathbf{X}_{t-1};\xi_{t})$. This modification is crucial as it is designed to directly cancel the variance introduced by the mini-batch $\xi_{t}$ Cutkosky2019MomentumBasedVR; Huang2021SUPERADAMFA; Yuan2024MARSUT. However, this theoretical benefit comes at the cost of requiring two gradient evaluations per step. MVR1 can be formally understood as a practical approximation of MVR2. The difference between their update rules is a single noise term, $\Delta^{\text{Noise}}_{t-1}=\nabla f(\mathbf{X}_{t-1};\xi_{t})-\nabla f(\mathbf{X}_{t-1};\xi_{t-1})$. Under the standard assumption of bounded variance ($\mathbb{E}_{\xi}[\|\nabla f(\mathbf{X};\xi)-\nabla f(\mathbf{X})\|_{F}^{2}]\leq\sigma^{2}$), the variance of this noise is well-controlled, satisfying $\mathbb{E}[\|\Delta^{\text{Noise}}_{t-1}\|_{F}^{2}]\leq 2\sigma^{2}$. While MVR1 is often sufficient in practice, this structural difference leads to fundamentally different theoretical guarantees. The principled variance cancellation in MVR2 allows our algorithm to achieve a superior iteration complexity of $\tilde{\mathcal{O}}(T^{-1/3})$, as we will formally establish in Theorem 3.2.